The ratio of the electric force between two electron to the gravitational force between them is of the order of

Grade:12

2 Answers

Vikas TU
6 years ago
Dear Student,
Fgravitational=Gm1m2/r2
=Gm2/r2

Felectrostatic=kq1q2/r2
=kq2/r2

Ratio= Felectrostatic/Fgravitational = (kq2/r2)/(Gm2/r2)
=kq2/Gm2
=0.379x1032.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)

 
Ashutosh
6 years ago
The ratio of electric force and gravitational force between two electrons is given below => E.F/G.F = 4.17 × 10^42